at a local clinic there are 8 men, 5 women,and 3 children in the waiting room.If three patients are randomly selected, find the probability that there is at least 1 child among them.
Answer by stanbon(75887)

at a local clinic there are 8 men, 5 women,and 3 children in the waiting room.If three patients are randomly selected, find the probability that there is at least 1 child among them.
---
P(at least one child) = 1 - P(no child)
P(x>=1) = 1 - 3C0(3/16)^0*(13/16)^3 = 1-(13/16)^3 = 0.4636
====================
cheers,
Stan H.